Let y=5x2+7x+3

Find the differential dy when x=2 and dx=0.4

Find the differential dy when x=2 and dx=0.8
Answer #1

Solution given that y=502+7x+3 Power rale dy = dx e or (29)=a.294 die (58²7 7x+3) dy = (10x+7) dx 0 when x=2, and da=0.4 dy -